OpenStudy (anonymous):

write an equation in standard form for the line that has undefined slope and passes through (-6,4).

OpenStudy (anonymous):

y=mx+c. Plug ur values.

OpenStudy (anonymous):

phil any probs?

OpenStudy (anonymous):

uhh....what values?

OpenStudy (anonymous):

x=-6 and y=4

OpenStudy (anonymous):

plug them to y=mx+c.You hav an equation in m and c. thats ur answer!

OpenStudy (anonymous):

so 4=m(-6)+c?

OpenStudy (anonymous):

yep..

OpenStudy (anonymous):

uhh............okay?.....ill try....

OpenStudy (anonymous):

any more probs? ask me here..

OpenStudy (anonymous):

if u didnt understand clearly...ask me thru!

OpenStudy (anonymous):

ok uhh.......idk.........how do u find the slope?

OpenStudy (anonymous):

i think u know whats a slope...its tangent of the angle made by the line with the positive x-axis...

OpenStudy (anonymous):

in the equation of a line of the from y=mx+c x and y are the coordinates of any point on the line...m is the slope...and c is the y-intercept(i.e. the point where the line cuts the y-axis)..clear?

OpenStudy (anonymous):

so we need 2 things to determine a line m and c..

OpenStudy (anonymous):

theres no Y intercept in an undefined

OpenStudy (anonymous):

well we are given only one information ..the coordinate of a point..so we have undetermined constants..

OpenStudy (anonymous):

yes...so we dont know c we keep it in the answer :)

OpenStudy (anonymous):

ask if u are not comfortable! i am here all along...:)

OpenStudy (anonymous):

i REALLY dont get what ur trying to tell me.

OpenStudy (mattfeury):

it will be x = -6, which is a straight vertical line that passes through that point. vertical lines have undefined slope since slope = rise / run (run is 0).

OpenStudy (anonymous):

k. if by undefined u mean a slope of tan90 then the line is perpendicular to x-axis..

OpenStudy (anonymous):

then mattfeury's answer is correct

OpenStudy (anonymous):

ok................................................................................................OH NOW GET IT

OpenStudy (mattfeury):

if we helped, click the good answer button to give us medals. sorry to snipe while yre helping, saubhik :p